OpenStudy (anonymous):

) linear inequality graph Which is an equation of the given line in standard form? A. –11x + 6y = –13 B. –6x +11y = 13 C. –6x + 11y = –13 D. –6x + 7y = –17

OpenStudy (anonymous):

(1 1/2,2) x1 = 1 1/2 or 3/2 and y1 = 2 (-4,-1) x2 = -4 and y2 = -1 find the slope.. slope(m) = (y2 - y1) / (x2 - x1) m = (-1-2) / (-4 - 3/2) m = -3/(-8/2 - 3/2) m = -3/(-11/2) m = -3 * -2/11 m = 6/11 y = mx + b -1 = 6/11(-4) + b -1 = - 24/11 + b -1 + 24/11 = b - 11/11 + 24/11 = b 13/11 = b y = 6/11x + 13/11 (multiply by 11) 11y = 6x + 13 -6x + 11y = 13 <----yes..answer is B

OpenStudy (anonymous):

thank u so much can u help me with a few more

OpenStudy (anonymous):

) linear inequality graph Which is the equation of the given line? A. x = –3 B. x = 3 C. y = –3 D. y = –3x

OpenStudy (anonymous):

@mathwiz916

OpenStudy (anonymous):

@just_one_last_goodbye